OpenStudy (anonymous):

which of the following is the equation of the exponential function y=Ce^kx that passes through (0,5) and (1,5e) a)y=5e^kx b)y=ce^5x c)y=5e^5x d)y=5e^-x e)none

OpenStudy (anonymous):

i have so far that c=5

OpenStudy (anonymous):

for the other one i plugged it in and so far i have 5e=5e^k how do i simplify it

OpenStudy (agent0smith):

C=5 is correct. For this one... put a 1 on the e, since e is really e^1\[\\[\Large 5 \times e^1=5 \times e^k\]\] then see if you can simplify

OpenStudy (agent0smith):

\[\Large 5 \times e^1 =5 \times e^k\]

OpenStudy (anonymous):

so then would my answer be that y=5e^kx

OpenStudy (agent0smith):

You'll need to simplify \[\Large 5 \times e^1 =5 \times e^k\] you can get rid of those 5's

OpenStudy (anonymous):

but then i would be left with e^1 =e^k

OpenStudy (agent0smith):

Then you can find k :)

OpenStudy (agent0smith):

\[\Large \Large e^1 = e^k \]then k must be...

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

so the answer would be none right since the final would have to be y=5e^x

OpenStudy (agent0smith):

Yep, looks that way.

OpenStudy (anonymous):

thanks